OpenStudy (anonymous):

A ranger spots a forest fire while on a 30-meter observation tower. The angle of depression from the tower to the fire is 20 degrees to the nearest meter how far is the fire from the base of the tower

OpenStudy (anonymous):

hmm let me go get my notes

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

so this is your equation now\[\tan (20) = 30/x\]

OpenStudy (anonymous):

so its .36 and now you want ot multiply both sides by x \[.36(x)=30/x *x \] this crosses out the x on the right.. now you're going to want to put the .36 on the right and leave x by its own

OpenStudy (travisbrown372):

tan 27 = 40 / x x = 40 / tan 27 x = 78.5 or 79 meters So the answer is C. The fire is 79 meters away from the observation tower.

OpenStudy (anonymous):

so you're going to divide both sides by .36 \[.36x/.36 = 30/.36\] \[x= 30/.36\]

OpenStudy (anonymous):

all you have to do is 30/.36 which turns out to be 83.3 so it will 83.3 feet

OpenStudy (anonymous):

Thank you for explain the problem I've been stuck on it for awhile

OpenStudy (anonymous):

i can find you a website ifyou'd like